The Rivoli Company has no debt outstanding, and its financial position is given by the following data:

Expected EBIT $600,000
Growth rate in EBIT, gL 0 %
Cost of equity, rs 10 %
Shares outstanding, no 100,000
Tax rate, T (federal-plus-state) 25 %
  1. What is Rivoli's intrinsic value of operations (i.e., its unlevered value)? Round your answer to the nearest dollar.

    What is its intrinsic stock price?

  2. Its earnings per share? Round your answers to the nearest cent.

    Intrinsic stock price: $   
Answer #1

a.

Let, all the earnings are distributed as dividend

Dividend = EBIT*(1 - Tax Rate) / No. of Shares

= 600,000 * (1 - 0.25) / 100,000

Dividend = $4.5

As there is no growth,

Stock Price = Dividend / WACC (As, there is no debt so WACC = Cost of Equity)

= $4.5 / 10%

Stock Price = $45

b.

EPS = EBIT * (1 - Tax Rate) / No. of Shares

= 600,000 * (1 - 0.25) / 100,000

EPS = $4.5